A half ton truck carrying sand , to capacity breaks down. The owner sends quarter ton truck to load the sand and deliver it to the client. How would you calculate the number of quarter ton trucks needed to curry the sand?

a. 1/2 × 2

b. 1/4 + 1/2

C. 1/4 × 2/1

D. 1/2 + 1/4



1
Expert's answer
2021-05-07T12:30:01-0400

C. 1/4 × 2/1


We need two quarter ton trucks to load the sand from one half ton truck.

"2 \times \frac{1}{4} = \frac{2}{4} = \frac{1}{2}"